OpenStudy (anonymous):

A 1,000 kilogram car travelling with the velocity of +20. m/s decelerates uniformly at -5.0 m/s^2 until it comes to rest. What is the total distance the car travels as it decelerates to rest?

OpenStudy (anonymous):

why not try the formula: \[v^2=u^2+2as\] where v=0 m/s u=20m/s a=-5m/s and solve for s

OpenStudy (anonymous):

It says here the answer (the distance) is 40 meters. I just can't seem to put that into place. Anyway, I'll try to find the 's' just like what you've suggested. Thanks. :)

OpenStudy (anonymous):

Ignore the weight of the car as you do not need it to calculate the distance travelled.

OpenStudy (anonymous):

@SweetsAndEverything do you need me to explain further?

OpenStudy (anonymous):

if you don't mind. :)

OpenStudy (anonymous):

the equation is: v^2=u^2+2as substitute the values in: v=final velocity (0 because you've stopped) u=initial velocity (20 because that is the inital speed) a=acceleration (-5 because your stopping) s=distance travelled